# The Bisognano-Wichmann property for non-unitary Wightman conformal field theories

### v1

@@ 5 Haag duality for Wightman CFTs with invariant Hermitian forms
which establishes condition (i) of the Lemma. Condition (ii) is straightforward, so the Lemma says that $\theta{\mathcal{P}}(I_{+})^{\prime}\theta\subset{\mathcal{P}}(I_{+})^{\prime\prime}$, as was required to complete the proof in case (a).
Assumption (b) implies (a) by Corollary 5.13, so we move on to assumption (c). As above, we need to show $\theta{\mathcal{Q}}(I_{+})\theta\subset{\mathcal{Q}}(I_{+})^{\prime}$. We will invoke Lemma 5.17 with $A={\mathcal{Q}}(I_{+})={\mathcal{P}}(I_{+})^{\prime\prime}$ and $B={\mathcal{P}}(I_{-})$. We first verify condition (i) of the Lemma. Let $X,Y\in{\mathcal{Q}}(I_{+})_{sa}$. By assumption (c), we may choose nets $X_{n}$ and $Y_{m}$ in ${\mathcal{P}}(I_{+})$ such that $X_{n}\to X$ and $Y_{m}\to Y$ in the weak operator topology. Replacing $X_{n}$ by $\tfrac{1}{2}(X_{n}+X_{n}^{*})$, and similarly for $Y_{n}$, we may assume that $X_{n}=X_{n}^{*}$ and $Y_{n}=Y_{n}^{*}$. Then $\theta X_{n}\Omega\in{\mathcal{K}}(I_{-})$ and $Y_{n}\Omega\in{\mathcal{K}}(I_{+})$, so we have
$\langle\theta X_{n}\Omega,Y_{m}\Omega\rangle\in{\mathbb{R}}.$
Taking limits first in $n$ and then in $m$ yields $\langle\theta X\Omega,Y\Omega\rangle\in{\mathbb{R}}$, establishing (i). Condition (ii) is straightforward, and we conclude that $\theta{\mathcal{Q}}(I_{+})\theta\subset{\mathcal{Q}}(I_{+})^{\prime}$, completing the proof. ∎
@@ 6 Applications for unitary CFTs
Recall that we considered the nets of algebras ${\mathcal{Q}}(I)={\mathcal{P}}(I)^{\prime\prime}$, where the commutant is taken in $\operatorname{End}_{*}({\mathcal{D}})$. This is a local Möbius covariant net of *-algebras, and we obtained in Section 5 several conditions which implied Haag duality for this net of algebras. As a result, we can show the following.
Corollary 6.2.
Let $({\mathcal{F}},{\mathcal{D}},U,\Omega)$ be a unitary Wightman CFT. Let ${\mathcal{P}}(I)\subset\operatorname{End}_{*}({\mathcal{D}})$ be the unital algebra generated by operators $\varphi(f)$ with $\varphi\in{\mathcal{F}}$ and $\operatorname{supp}f\subset I$, and let ${\mathcal{Q}}(I)={\mathcal{P}}(I)^{\prime\prime}$. Then ${\mathcal{Q}}(I)$ is a Haag dual Möbius covariant net of algebras.
Proof.
The net ${\mathcal{K}}_{w}(I)$ is Haag dual by Proposition 6.1, and so ${\mathcal{Q}}$ is Haag dual by Proposition 5.18. ∎
Recall from Section 2.1 that given a standard subspace $K$ of a Hilbert space ${\mathcal{H}}$, we have a corresponding conjugate linear closed involution $S_{K}$ with polar decomposition $S_{K}=J\Delta^{1/2}$. Given a Móbius covariant net of standard subspaces ${\mathcal{K}}(I)\subset{\mathcal{H}}$, covariant with respect to the unitary representation $U$ of $\operatorname{M\ddot{o}b}$ on ${\mathcal{H}}$, then the Bisognano-Wichmann property for ${\mathcal{K}}$ [Lon08a, Thm. 3.3.1] states that the modular automorphisms $\Delta_{\pm}^{it}$ for ${\mathcal{K}}(I_{\pm})$ are given by
@@
Let $({\mathcal{F}},{\mathcal{D}},U,\Omega)$ be a unitary Wightman CFT with PCT operator $\theta$, and let ${\mathcal{H}}$ be the Hilbert space completion of ${\mathcal{D}}$. Let $\hat{\mathcal{K}}_{n}(I)\subset{\mathcal{H}}$ be the corresponding net of standard subspaces defined as the closure of ${\mathcal{P}}(I)_{sa}\Omega$. Let $V_{\pm i\pi}$ be the densely defined operators on ${\mathcal{D}}$ defined in Definition 4.1. Let
±=${\mathcal{E}}_{\pm}=\{\Phi\in D(\Delta^{1/2}_{\pm})\,:\,\Phi\in{\mathcal{D}}\text{ and }\Delta_{\pm}^{1/2}\Phi\in{\mathcal{D}}\}.$
@@
$F_{\Phi^{\prime}}(t)=\langle V_{t}\Phi,\Phi^{\prime}\rangle,\qquad F_{\Phi^{\prime}}(t+i\pi)=\langle V_{t}\Delta^{1/2}\Phi,\Phi^{\prime}\rangle.$
Hence if $\Phi,\Delta^{1/2}_{+}\Phi\in{\mathcal{D}}$, then it follows that $\Phi\in D(V_{i\pi})$ and $V_{i\pi}\Phi=\Delta^{1/2}\Phi$.
It remains to check that ±${\mathcal{E}}_{\pm}=D(V_{\pm i\pi})$, and since we have shown ±${\mathcal{E}}_{\pm}\subset D(V_{\pm i\pi})$, it only remains to check the opposite inclusion. As in Lemma 5.12, let ±${\mathcal{L}}_{\pm}\subset{\mathcal{D}}$ be the standard subspace corresponding to the involution $S_{\pm}=\theta V_{\pm i\pi}$, so that $D(V_{\pm i\pi})={\mathcal{L}}_{\pm}+i{\mathcal{L}}_{\pm}$. By the same lemma, ${\mathcal{L}}_{\pm}^{\prime}={\mathcal{L}}_{\mp}$ and ${\mathcal{K}}_{w}(I_{\pm})\subset{\mathcal{L}}_{\pm}$. By Proposition 6.1, the net ${\mathcal{K}}_{w}(I_{\pm})$ is Haag dual, so taking complements yields ${\mathcal{L}}_{\mp}\subset{\mathcal{K}}_{w}(I_{\mp})$. Thus we have shown ±${\mathcal{L}}_{\pm}={\mathcal{K}}_{w}(I_{\pm})$. By Proposition 6.1 we have ${\mathcal{K}}_{w}(I_{\pm})={\mathcal{K}}_{n}(I_{\pm})=\hat{\mathcal{K}}_{n}(I_{\pm})\cap{\mathcal{D}}$. Putting it all together, we conclude that ±${\mathcal{L}}_{\pm}=\hat{\mathcal{K}}_{n}(I_{\pm})\cap{\mathcal{D}}$. We thus have:
@@
For analytic details, see [RTT22, §2.4] (and also [CKLW18, CRTT24]). We now let ${\mathcal{F}}$ be the Wightman CFT associated with the generating set of all quasiprimary vectors $v$, i.e.
=${\mathcal{F}}=\{Y(v,\cdot)\,:\,v\text{ quasiprimary}\},$
@@
${\mathcal{H}}_{\mathcal{C}}=\overline{\left(\bigvee_{I\subset S^{1}}{\mathcal{C}}(I)\right)\Omega},$
where $\bigvee_{I}{\mathcal{C}}^{\mathcal{H}}(I)$ is the von Neumann algebra generated by the algebras ${\mathcal{C}}(I)$. Let $p_{\mathcal{C}}$ be the projection of ${\mathcal{H}}$ onto ${\mathcal{H}}_{\mathcal{C}}$, and let ${\mathcal{C}}(I)={\mathcal{C}}^{\mathcal{H}}(I)p_{\mathcal{C}}$, the cutdown of the net of algebras ${\mathcal{C}}^{\mathcal{H}}$ to its vacuum Hilbert space ${\mathcal{H}}_{\mathcal{C}}$.
@@
The real subspaces $\overline{{\mathcal{C}}(I)_{sa}\Omega}$ are a Möbius-covariant net of standard subspaces of the Hilbert space ${\mathcal{H}}_{\mathcal{C}}$. The modular conjugation corresponding to the upper semicircle is $\theta|_{{\mathcal{H}}_{\mathcal{C}}}$.
@@
We verify that these subspaces satisfy Definition 2.3. Isotony of the net is clear, as is the spectrum condition. The algebras ${\mathcal{C}}^{\mathcal{H}}(I)$ are Möbius covariant with respect to the given unitary representation on ${\mathcal{H}}$. It follows that ${\mathcal{H}}_{\mathcal{C}}$ is invariant under the Möbius group, and that the algebras ${\mathcal{C}}(I)$ are Möbius covariant, and indeed that the subspaces $\overline{{\mathcal{C}}(I)_{sa}\Omega}$ are as well. It remains to check the cyclicity condition. Let ~$\tilde{\mathcal{H}}_{\mathcal{C}}\subset{\mathcal{H}}_{\mathcal{C}}$ be the closed span of the subspaces ${\mathcal{C}}(I)\Omega$. Then evidently the given subspaces form a Möbius-covariant net of subspaces of ~$\tilde{\mathcal{H}}_{\mathcal{C}}$. Then by the Reeh-Schlieder property [Lon08a, Thm. 3.2.1] we have ~$\tilde{\mathcal{H}}_{\mathcal{C}}=\overline{{\mathcal{C}}(I)\Omega}$ for each interval $I$. Hence ~$\tilde{\mathcal{H}}_{\mathcal{C}}$ is invariant under each algebra ${\mathcal{C}}(I)$. It follows that ~$\tilde{\mathcal{H}}_{\mathcal{C}}$ is invariant under $\bigvee_{I}{\mathcal{C}}(I)$, from which it follows that ${\mathcal{H}}_{\mathcal{C}}\subset\tilde{\mathcal{H}}_{\mathcal{C}}$. Hence ${\mathcal{H}}_{\mathcal{C}}=\tilde{\mathcal{H}}_{\mathcal{C}}$, which completes the proof of the locality axiom. ∎
@@
Let ${\mathcal{V}}$ be a unitary Möbius vertex algebra with PCT operator $\theta$, let ${\mathcal{C}}^{\mathcal{H}}(I)={\mathcal{A}}_{\mathcal{V}}(I^{\prime})^{\prime}$ be the dual net of von Neumann algebras, and let ${\mathcal{C}}(I)$ be the cutdown of ${\mathcal{C}}^{\mathcal{H}}(I)$ to the vacuum Hilbert space ${\mathcal{H}}_{\mathcal{C}}$, as described above. Then ${\mathcal{C}}$ is a a Möbius-covariant Haag-Kastler net on ${\mathcal{H}}_{\mathcal{C}}$. The modular conjugation associated with ${\mathcal{C}}(I_{+})$ and $\Omega$ is $\theta|_{{\mathcal{H}}_{\mathcal{C}}}$.
@@
The only axiom which needs to be checked is locality. This will follow from the assertion regarding the modular conjugation, which we now verify. Let ${\mathcal{N}}=\overline{{\mathcal{C}}(I_{+})_{sa}\Omega}$, which is a standard subspace of ${\mathcal{H}}_{\mathcal{C}}$ by Lemma 6.5. We note that it follows that $\Omega$ is cyclic and separating for ${\mathcal{C}}(I_{+})$. By definition ${\mathcal{C}}^{\mathcal{H}}(I_{+})$ commutes with ${\mathcal{P}}(I_{-})$, and therefore ${\mathcal{C}}(I_{+})_{sa}\Omega\subset\hat{\mathcal{K}}_{n}(I_{-})^{\prime}=\hat{\mathcal{K}}_{n}(I_{+})$. It follows that ${\mathcal{N}}\subset\hat{\mathcal{K}}_{n}(I_{+})$.
From the theory of Möbius covariant nets of subspaces we know that the modular automorphism group of $\hat{\mathcal{K}}_{n}(I_{+})$ is $\Delta_{+}^{it}=U(v_{-2\pi t})$ [Lon08a, Thm. 3.3.1]. As ${\mathcal{N}}$ is invariant under $\Delta_{+}^{it}$, the modular automorphism group for ${\mathcal{N}}$ is $\Delta_{+}^{it}|_{{\mathcal{H}}_{\mathcal{N}}}$ [Lon08a, Cor. 2.1.8]. Let $S_{\mathcal{N}}$ be the involution on ${\mathcal{H}}_{\mathcal{N}}$ corresponding to ${\mathcal{N}}$, and let $S_{\mathcal{N}}=J_{\mathcal{N}}\Delta_{\mathcal{N}}^{1/2}$ be the polar decomposition. For $\xi\in D(\Delta_{\mathcal{N}}^{1/2})$, the map $\Delta_{\mathcal{N}}^{it}\xi$ extends to a continuous function on the strip $\{-\tfrac{1}{2}\leq\operatorname{Im}(\tau)\leq 0\}$, holomorphic on the interior, mapping $\tau=-i/2$ to $\Delta^{1/2}_{\mathcal{N}}\xi$. As the same holds for $\Delta_{+}^{it}$, we have $\Delta^{1/2}_{\mathcal{N}}\xi=\Delta^{1/2}_{+}\xi$ for $\xi\in{\mathcal{N}}+i{\mathcal{N}}$, and in particular for $\xi$ of the form $x\Omega$ with $x\in{\mathcal{C}}(I_{+})$. Thus if $x\in{\mathcal{C}}(I_{+})$ we have
@@
with the last equality by Theorem 6.3. Hence by Reeh-Schlieder (which in this case follows from Lemma 6.5) we have $J_{\mathcal{N}}=\theta|_{{\mathcal{H}}_{\mathcal{N}}}$, proving our claim about the modular conjugation for ${\mathcal{C}}(I_{+})$.
We now complete the proof of locality, for which it suffices to show that ${\mathcal{C}}(I_{+})$ and ${\mathcal{C}}(I_{-})$ commute. By the main theorem of Tomita-Takesaki theory, we have $\theta{\mathcal{C}}(I_{+})\theta={\mathcal{C}}(I_{+})^{\prime}$. On the other hand, since $\theta{\mathcal{A}}_{\mathcal{V}}(I_{\pm})\theta={\mathcal{A}}_{\mathcal{V}}(I_{\mp})$, it follows that $\theta{\mathcal{C}}^{\mathcal{H}}(I_{+})\theta={\mathcal{C}}^{\mathcal{H}}(I_{-})$. Since $\theta$ leaves ${\mathcal{H}}_{\mathcal{N}}$ invariant we have $\theta{\mathcal{C}}(I_{+})\theta={\mathcal{C}}(I_{-})$. Hence ${\mathcal{C}}(I_{+})^{\prime}={\mathcal{C}}(I_{-})$, completing the proof of locality. ∎
@@
    $\Omega$ is separating for ${\mathcal{A}}_{V}(I)$, for some (or any) interval $I$
@@
If ${\mathcal{V}}$ is AQFT-local, then ${\mathcal{A}}_{\mathcal{V}}(I)$ is a Möbius-covariant Haag-Kastler net, and $\Omega$ is separating for ${\mathcal{A}}_{\mathcal{V}}(I)$ by the Reeh-Schlieder property. Conversely, if $\Omega$ is separating, then it is cyclic for ${\mathcal{C}}^{\mathcal{H}}(I)={\mathcal{A}}_{\mathcal{V}}(I^{\prime})^{\prime}$, and ${\mathcal{H}}_{\mathcal{C}}={\mathcal{H}}$. Hence by Theorem 6.6 , ${\mathcal{C}}(I)$ is a Möbius-covariant Haag Kastler net on ${\mathcal{H}}$, and by Haag duality for that net we have ${\mathcal{C}}(I)={\mathcal{A}}_{\mathcal{V}}(I)$, completing the proof. The computation of the modular conjugation follows from Theorem 6.6 as well. ∎
We note that the claim about the modular conjugation could have alternatively been proven directly from [CKLW18, Thm. B.6], under the assumption of energy bounds.

### v2

@@ 5 Haag duality for Wightman CFTs with invariant Hermitian forms
which establishes condition (i) of the Lemma. Condition (ii) is straightforward, so the Lemma says that $\theta{\mathcal{P}}(I_{+})^{\prime}\theta\subset{\mathcal{P}}(I_{+})^{\prime\prime}$, which completes the proof. ∎
@@ 6 Applications for unitary CFTs
Recall from Section 2.1 that given a standard subspace $K$ of a Hilbert space $\mathcal{H}$}, we have a corresponding conjugate linear closed involution $S_{K}$ with polar decomposition $S_{K}=J\Delta^{1/2}$. Given a Möbius covariant net of standard subspaces ${\mathcal{K}}(I)\subset{\mathcal{H}}$, covariant with respect to the unitary representation $U$ of $\operatorname{M\ddot{o}b}$ on $\mathcal{H}$}, then the Bisognano-Wichmann property for $\mathcal{K}$} [Lon08a, Thm. 3.3.1] states that the modular automorphisms $\Delta_{\pm}^{it}$ for ${\mathcal{K}}(I_{\pm})$ are given by
@@
Let $({\mathcal{F}},{\mathcal{D}},U,\Omega)$ be a unitary Wightman CFT with PCT operator $\theta$, and let $\mathcal{H}$} be the Hilbert space completion of $\mathcal{D}$}. Let $\hat{\mathcal{K}}_{n}(I)\subset{\mathcal{H}}$ be the corresponding net of standard subspaces defined as the norm closure of ${\mathcal{P}}(I)_{sa}\Omega$ in $\mathcal{H}$}. Let $V_{\pm i\pi}$ be the densely defined operators on $\mathcal{D}$} defined in Definition 4.1. Let
${\mathcal{E}}_{\pm}=\{\Phi\in D(\Delta^{1/2}_{\pm})\,:\,\Phi\in{\mathcal{D}}\text{ and }\Delta_{\pm}^{1/2}\Phi\in{\mathcal{D}}\}.$
@@
$F_{\Phi^{\prime}}(t)=\langle V_{t}\Phi,\Phi^{\prime}\rangle,\qquad F_{\Phi^{\prime}}(t+i\pi)=\langle V_{t}\Delta_{+}^{1/2}\Phi,\Phi^{\prime}\rangle.$
Hence if $\Phi,\Delta^{1/2}_{+}\Phi\in{\mathcal{D}}$, then it follows that $\Phi\in D(V_{i\pi})$ and $V_{i\pi}\Phi=\Delta_{+}^{1/2}\Phi$.
It remains to check that ${\mathcal{E}}_{\pm}=D(V_{\pm i\pi})$, and since we have shown ${\mathcal{E}}_{\pm}\subset D(V_{\pm i\pi})$, it only remains to check the opposite inclusion. As in Lemma 5.12, let ${\mathcal{L}}_{\pm}\subset{\mathcal{D}}$ be the standard subspace corresponding to the involution $S_{\pm}=\theta V_{\pm i\pi}$, so that $D(V_{\pm i\pi})={\mathcal{L}}_{\pm}+i{\mathcal{L}}_{\pm}$. By the same lemma, $\mathcal{L}$$_{\pm}^{\prime}$$\mathcal{L}$$_{\mp}$ and ${\mathcal{K}}_{w}(I_{\pm})\subset{\mathcal{L}}_{\pm}$. By Proposition 6.1, the net ${\mathcal{K}}_{w}(I_{\pm})$ is Haag dual, so taking complements yields ${\mathcal{L}}_{\mp}\subset{\mathcal{K}}_{w}(I_{\mp})$. Thus we have shown ${\mathcal{L}}_{\pm}={\mathcal{K}}_{w}(I_{\pm})$. By Proposition 6.1 we have ${\mathcal{K}}_{w}(I_{\pm})={\mathcal{K}}_{n}(I_{\pm})=\hat{\mathcal{K}}_{n}(I_{\pm})\cap{\mathcal{D}}$. Putting it all together, we conclude that ${\mathcal{L}}_{\pm}=\hat{\mathcal{K}}_{n}(I_{\pm})\cap{\mathcal{D}}$. We thus have:
@@
For analytic details, see [RTT22, §2.4] (and also [CKLW18] for a treatment in the presence of energy bounds). We now let $\mathcal{F}$} be the Wightman CFT associated with the generating set of all quasiprimary vectors $v$, i.e.
${\mathcal{F}}=\{Y(v,\cdot)\,:\,v\text{ quasiprimary}\},$
@@
${\mathcal{H}}_{\mathcal{C}}=\overline{\left(\bigvee_{I\subset S^{1}}{\mathcal{C}}^{\mathcal{H}}(I)\right)\Omega},$ (6.2)
where $\bigvee_{I}{\mathcal{C}}^{\mathcal{H}}(I)$ is the von Neumann algebra generated by the algebras ${\mathcal{C}}^{\mathcal{H}}(I)$. Let $p_{\mathcal{C}}$ be the projection of $\mathcal{H}$} onto ${\mathcal{H}}_{\mathcal{C}}$, and let ${\mathcal{C}}(I)={\mathcal{C}}^{\mathcal{H}}(I)p_{\mathcal{C}}$, the cutdown of the net of algebras $\mathcal{C}$$^{\mathcal{H}}$ to its vacuum Hilbert space ${\mathcal{H}}_{\mathcal{C}}$.
@@
The real subspaces $\overline{{\mathcal{C}}(I)_{sa}\Omega}$ are a Möbius-covariant net of standard subspaces of the Hilbert space ${\mathcal{H}}_{\mathcal{C}}$.
@@
We verify that these subspaces satisfy Definition 2.3. Isotony of the net is clear, as is the spectrum condition. The algebras ${\mathcal{C}}^{\mathcal{H}}(I)$ are Möbius covariant with respect to the given unitary representation on $\mathcal{H}$}. It follows that ${\mathcal{H}}_{\mathcal{C}}$ is invariant under the Möbius group, and that the algebras ${\mathcal{C}}(I)$ are Möbius covariant, and indeed that the subspaces $\overline{{\mathcal{C}}(I)_{sa}\Omega}$ are as well.
We next check locality. By definition ${\mathcal{C}}^{\mathcal{H}}(I)$ commutes with ${\mathcal{P}}(I^{\prime})$, and therefore if $x\in{\mathcal{C}}^{\mathcal{H}}(I)_{sa}$ and $Y\in{\mathcal{P}}(I^{\prime})_{sa}$ we have $\langle x\Omega,Y\Omega\rangle\in{\mathbb{R}}$. Recall that $\hat{\mathcal{K}}_{n}(I^{\prime})$ is the norm closure of ${\mathcal{P}}(I^{\prime})_{sa}\Omega$ in $\mathcal{H}$}, and observe that it now follows that
${\mathcal{C}}(I)_{sa}\Omega={\mathcal{C}}^{\mathcal{H}}(I)_{sa}\Omega\subset\hat{\mathcal{K}}_{n}(I^{\prime})^{\prime}=\hat{\mathcal{K}}_{n}(I),$
with the last equality by Proposition 6.1. The locality of the subspaces $\overline{{\mathcal{C}}(I)_{sa}\Omega}$ now follows from the locality of the net $\hat{\mathcal{K}}_{n}(I)$.
It remains to check the cyclicity condition. Let $\tilde{\mathcal{H}}_{\mathcal{C}}\subset{\mathcal{H}}_{\mathcal{C}}$ be the closed span of the subspaces ${\mathcal{C}}(I)\Omega$. Then evidently the given subspaces form a Möbius-covariant net of subspaces of $\tilde{\mathcal{H}}_{\mathcal{C}}$. Then by the Reeh-Schlieder property [Lon08a, Thm. 3.2.1] we have $\tilde{\mathcal{H}}_{\mathcal{C}}=\overline{{\mathcal{C}}(I)\Omega}$ for each interval $I$. Hence $\tilde{\mathcal{H}}_{\mathcal{C}}$ is invariant under each algebra ${\mathcal{C}}(I)$. It follows that $\tilde{\mathcal{H}}_{\mathcal{C}}$ is invariant under $\bigvee_{I}{\mathcal{C}}(I)$, from which it follows that ${\mathcal{H}}_{\mathcal{C}}\subset\tilde{\mathcal{H}}_{\mathcal{C}}$. Hence ${\mathcal{H}}_{\mathcal{C}}=\tilde{\mathcal{H}}_{\mathcal{C}}$, which completes the proof of the cyclicity axiom. ∎
@@
Let $\mathcal{V}$} be a unitary Möbius vertex algebra with PCT operator $\theta$, let ${\mathcal{C}}^{\mathcal{H}}(I)={\mathcal{A}}_{\mathcal{V}}(I^{\prime})^{\prime}$ be the dual net of von Neumann algebras, and let ${\mathcal{C}}(I)$ be the cutdown of ${\mathcal{C}}^{\mathcal{H}}(I)$ to the vacuum Hilbert space ${\mathcal{H}}_{\mathcal{C}}$, as described above. Then $\mathcal{C}$} is a Möbius-covariant Haag-Kastler net on ${\mathcal{H}}_{\mathcal{C}}$. The modular conjugation associated with ${\mathcal{C}}(I_{+})$ and $\Omega$ is $\theta|_{{\mathcal{H}}_{\mathcal{C}}}$.
@@
The only axiom which needs to be checked is locality. This will follow from the assertion regarding the modular conjugation, which we now verify. Let ${\mathcal{N}}=\overline{{\mathcal{C}}(I_{+})_{sa}\Omega}$, which is a standard subspace of ${\mathcal{H}}_{\mathcal{C}}$ by Lemma 6.4. Recall from the proof of Lemma 6.4 that ${\mathcal{N}}\subset\hat{\mathcal{K}}_{n}(I_{+})$. From the theory of Möbius covariant nets of subspaces we know that the modular automorphism group of $\hat{\mathcal{K}}_{n}(I_{+})$ is $\Delta_{+}^{it}=U(v_{-2\pi t})$ [Lon08a, Thm. 3.3.1]. As $\mathcal{N}$} is invariant under $\Delta_{+}^{it}$, the modular automorphism group for $\mathcal{N}$} is $\Delta_{+}^{it}|_{{\mathcal{H}}_{\mathcal{C}}}$ [Lon08a, Cor. 2.1.8]. Let $S_{\mathcal{N}}$ be the involution on ${\mathcal{H}}_{\mathcal{C}}$ corresponding to $\mathcal{N}$}, and let $S_{\mathcal{N}}=J_{\mathcal{N}}\Delta_{\mathcal{N}}^{1/2}$ be the polar decomposition. For $\xi\in D(\Delta_{\mathcal{N}}^{1/2})$, the map $\Delta_{\mathcal{N}}^{it}\xi$ extends to a continuous function on the strip $\{-\tfrac{1}{2}\leq\operatorname{Im}(\tau)\leq 0\}$, holomorphic on the interior, mapping $\tau=-i/2$ to $\Delta^{1/2}_{\mathcal{N}}\xi$. As the same holds for $\Delta_{+}^{it}$, we have $\Delta^{1/2}_{\mathcal{N}}\xi=\Delta^{1/2}_{+}\xi$ for $\xi\in{\mathcal{N}}+i{\mathcal{N}}$, and in particular for $\xi$ of the form $x\Omega$ with $x\in{\mathcal{C}}(I_{+})$. Thus if $x\in{\mathcal{C}}(I_{+})$ we have
@@
with the last equality by Theorem 6.2. Hence by Reeh-Schlieder (which in this case follows from Lemma 6.4) we have $J_{\mathcal{N}}=\theta|_{{\mathcal{H}}_{\mathcal{C}}}$, proving our claim about the modular conjugation for ${\mathcal{C}}(I_{+})$.
We now complete the proof of locality, for which it suffices to show that ${\mathcal{C}}(I_{+})$ and ${\mathcal{C}}(I_{-})$ commute. By the main theorem of Tomita-Takesaki theory, we have $\theta{\mathcal{C}}(I_{+})\theta={\mathcal{C}}(I_{+})^{\prime}$. On the other hand, since $\theta{\mathcal{A}}_{\mathcal{V}}(I_{\pm})\theta={\mathcal{A}}_{\mathcal{V}}(I_{\mp})$, it follows that $\theta{\mathcal{C}}^{\mathcal{H}}(I_{+})\theta={\mathcal{C}}^{\mathcal{H}}(I_{-})$. Since $\theta$ leaves ${\mathcal{H}}_{\mathcal{C}}$ invariant we have $\theta{\mathcal{C}}(I_{+})\theta={\mathcal{C}}(I_{-})$. Hence ${\mathcal{C}}(I_{+})^{\prime}={\mathcal{C}}(I_{-})$, completing the proof of locality. ∎
@@
    $\Omega$ is separating for ${\mathcal{A}}_{\mathcal{V}}(I)$, for some (or any) interval II
@@
If $\mathcal{V}$} is AQFT-local, then ${\mathcal{A}}_{\mathcal{V}}(I)$ is a Möbius-covariant Haag-Kastler net, and $\Omega$ is separating for ${\mathcal{A}}_{\mathcal{V}}(I)$ by the Reeh-Schlieder property. Conversely, if $\Omega$ is separating, then it is cyclic for ${\mathcal{C}}^{\mathcal{H}}(I)={\mathcal{A}}_{\mathcal{V}}(I^{\prime})^{\prime}$, and ${\mathcal{H}}_{\mathcal{C}}={\mathcal{H}}$. Hence by Theorem 6.5, ${\mathcal{C}}(I)$ is a Möbius-covariant Haag-Kastler net on $\mathcal{H}$}, and by Haag duality for that net [GF93, Thm. 2.19] we have ${\mathcal{C}}(I)={\mathcal{A}}_{\mathcal{V}}(I)$, completing the proof. The computation of the modular conjugation follows from Theorem 6.5 as well. ∎
We note that the claim about the modular conjugation could have alternatively been proven directly from [CKLW18, Thm. B.6], under the assumption of energy bounds. We also note that the proof of Corollary 6.6 also shows that $\mathcal{V}$} is AQFT-local when $\Omega$ is separating for $\bigcap_{I}{\mathcal{A}}_{\mathcal{V}}(I)$.